Two planes which are 3650 miles apart fly towards each other. Their speeds differ by 80 mph. If they pass each other in 5 hours what is the speed of each?
Found 3 solutions by ikleyn, MathTherapy, greenestamps:
Answer by ikleyn(52877)

.
Your equation for the speed of the slower plane is

5x + 5(x+80) = 3650.

==========================================>

10x + 400 = 3650  ====>  10x = 3650 - 400 = 3250 ====>  x = 325 mph.


Answer. Slower plane speed 325 mph.  Faster plane speed 325+80 = 405 mph.

Solved.


It is how simple is it.


---------------
See the lesson
- Travel and Distance problems
- Travel and Distance problems for two bodies moving in opposite directions
in this site.


Answer by MathTherapy(10556)


Two planes which are 3650 miles apart fly towards each other. Their speeds differ by 80 mph. If they pass each other in 5 hours what is the speed of each?
 


Answer by greenestamps(13209)


The distance of 3650 miles is covered by the two planes together in 5 hours, so the combined speeds of the two planes is 3650/5 = 730.

You can use formal algebra if you want to find the speed of each plane, knowing that their combined speed is 730 mph and that one plane's speed is 80 mph greater then the other's.

Or you can subtract the "extra" 80 mph of the faster plane to find that the two planes flying at the same speed would have a combined speed of 730-80=650 mph. Then half of that for the slow plane is 325 mph, making the speed of the faster plane 325+80=405 mph.